Suppose that during a test drive of two cars, one car travels 333 miles in the same time that the second car travels 270 miles. If the speed of one car is 14 miles per hour faster than the speed of the second car, find the speed of both cars.
Answer by richwmiller(17219)

y=x+14,
y*t=333,
x*t=270
t = 9/2=4.5, x = 60, y = 74